OpenStudy (naruko885):

What is the volume of a cylinder with a radius of 9 inches and a height of 2 inches? Use 3.14 for pi. Round your answer to the nearest hundredth. 500.21 cubic inches 508.68 cubic inches 550.45 cubic inches 561.75 cubic inches

OpenStudy (revlogic):

Volume = pi*r^2*h V = 3.14*9^2*2 V = 3.14*81*2 V = 508.68 cubic inches